# THOR: Tool-Integrated Hierarchical Optimization via RL for Mathematical Reasoning

### v2

@@ Abstract
Large Language Models (LLMs) have made remarkable progress in mathematical reasoning, but still continue to struggle with high-precision tasks like numerical computation and formal symbolic manipulation. Integrating external tools has emerged as a promising approach to bridge this gap. Despite recent advances, existing methods struggle with three key challenges: constructing tool-integrated reasoning data, performing fine-grained optimization, and enhancing inference. To overcome these limitations, we propose THOR (Tool-Integrated Hierarchical Optimization via RL). First, we introduce TIRGen, a multi-agent actor-critic-based pipeline for constructing high-quality datasets of tool-integrated reasoning paths, aligning with the policy and generalizing well across diverse models. Second, to perform fine-grained hierarchical optimization, we introduce an RL strategy that jointly optimizes for both episode-level problem solving and step-level code generation. This is motivated by our key insight that the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Finally, THOR incorporates a self-correction mechanism that leverages immediate tool feedback to dynamically revise erroneous reasoning paths during inference. Our approach demonstrates strong generalization across diverse models, performing effectively in both reasoning and non-reasoning models. It further achieves state-of-the-art performance for models of a similar scale on multiple mathematical benchmarks, while also delivering consistent improvements on code benchmarks. Our code will be publicly available at https://github.com/JingMog/THOR.
@@ 1 Introduction
Recent tool-free methods for enhancing LLMs’ mathematical reasoning can be broadly categorized into search-based methods (besta2024graph; zhang2025lessons; hu2025prm) and reinforcement learning (RL) paradigms (yu2025dapo; yue2025vapo). Despite notable progress, these approaches remain constrained by a fundamental weakness of LLMs. As probabilistic next-token predictors, they inherently struggle with high-precision tasks (chen2022program), such as numerical computation, equation solving, symbolic manipulation (pan2025enhancing), and formal proofs (lewkowycz2022solving). In contrast, programmatic reasoning excels in these domains. Therefore, integrating the semantic reasoning capabilities of LLMs with the precise and verifiable execution of external code-based tools provides a promising pathway to overcome these limitations.
@@
To address these challenges, we propose THOR, a tool-integrated framework designed to enhance the reasoning ability of LLMs. (1) For constructing TIR data, in order to efficiently generate policy-aligned TIR data, we propose TIRGen, an actor-critic-based data construction pipeline. The actor is responsible for generating natural language reasoning steps, while the critic evaluates whether steps can be transformed into executable code and interacts with an external executor to refine the reasoning. This iterative process yields a TIR dataset that is naturally aligned with the actor’s policy and broadly applicable across diverse models and tools. (2) For performing fine-grained optimization, we are motivated by the key insight that the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Our experiments later confirm this insight. Based on this, we introduce a hierarchical RL strategy that combines episode-level and step-level optimization. At the episode-level, we directly optimize for the correctness of the final answer. Concurrently, at the step-level, we apply fine-grained optimization to execution failure steps, specifically enhancing the model’s code generation ability. (3) For enhancing inference, we propose a self-correction mechanism that leverages immediate feedback from tools to dynamically revise its CoT during inference. When code invocation fails, it backtracks and explores alternative reasoning paths, thereby significantly enhancing the model’s reasoning robustness and overall performance.
@@ 2.2 TIRGen: TIR Data Generation Pipeline
Figure 2: The TIR data construction pipeline. In this pipeline, the Actor agent generates reasoning steps. The Critic agent identifies tool-executable steps and converts them into tool-augmented reasoning steps. After multi-stage filtering, we obtain the cold start dataset $\mathcal{D}_{SFT}$.
Existing methods for TIR highlight a significant need for high-quality training data. Most approaches rely on simply prompting external large models to synthesize TIR data for non-reasoning models (gou2023tora; li2025torl), but these approaches fail to generalize effectively to reasoning models such as DeepSeek-R1. Although START (li2025start) constructs long-CoT TIR data using a rule-based prompt-hint approach, the resulting trajectories often contain redundant code invocations. Therefore, existing TIR data construction methods face critical shortcomings, including style mismatches between the generated data and policy models, as well as poor scalability to reasoning models. To overcome these limitations, we introduce TIRGen, an automated TIR data synthesis pipeline based on an actor-critic framework, as illustrated in Figure 2.
In this framework, an Actor agent generates a natural language reasoning step with a maximum length of $L_{\text{step}}$. A Critic agent then evaluates this step to identify operations that can be more efficiently solved with code, such as numerical calculations or equation solving. Upon identifying such an operation, the Critic converts it into an executable Python code snippet while preserving the Actor’s original reasoning logic. The Critic then interacts with a code interpreter to obtain a precise execution result, which replaces the original operation and produces a code-augmented reasoning step. The Actor resumes reasoning from this revised step, and the iterative cycle continues until a complete solution is reached, as described in Algorithm 1. This design yields two key advantages:
  • Policy Alignment. Because the Critic observes only isolated reasoning steps without the full problem statement and final answer, the synthesized data directly reflects the Actor’s intrinsic reasoning preferences. This ensures that the data remains in-distribution, thereby mitigating the performance degradation commonly caused by training on out-of-distribution data (gudibande2023false; chen2024self).
  • Reduced Reliance on Large-scale Models. The Actor is responsible for high-level mathematical reasoning, while the Critic requires only basic instruction-following and code-generation skills. This division of labor decomposes the complex task into subproblems that are collaboratively solved by the two agents.
@@
1:Input: Actor model $\pi_{\text{actor}}$, Critic model $\pi_{\text{critic}}$, Dataset $\mathcal{D}_{q}$, Code interpreter sandbox $S$.
@@
6:   $r^{t}\sim\pi_{\text{actor}}(\cdot\mid\tau),\>|r^{t}|\leq L_{\text{step}}$ $\triangleright$ Actor $\pi_{\text{actor}}$ generates a reasoning step
@@
8:     $r^{t}_{\text{logic}}\leftarrow\text{ExtractLogic}_{\pi_{\text{critic}}}(r^{t})$ $\triangleright$ Step 1: Extract the pure reasoning part
9:     $a^{t}\leftarrow\text{ConvertToCode}_{\pi_{\text{critic}}}(r^{t},r^{t}_{\text{logic}})$ $\triangleright$ Step 2: Convert calculation part to code
@@ 2.3.2 Hierarchical Reinforcement Learning
where $s^{\prime}_{i}$ and $A^{\prime}_{i}$ denote the responses and advantages at the step-level.
The final training objective combines both episode-level and step-level losses:

### v3

@@ Abstract
Large Language Models (LLMs) have made remarkable progress in mathematical reasoning, but still continue to struggle with high-precision tasks like numerical computation and formal symbolic manipulation. Integrating external tools has emerged as a promising approach to bridge this gap. Despite recent advances, existing methods struggle with three key challenges: constructing tool-integrated reasoning data, performing fine-grained optimization, and enhancing inference. To overcome these limitations, we propose THOR (Tool-Integrated Hierarchical Optimization via RL). First, we introduce TIRGen, a multi-agent based pipeline for constructing high-quality datasets of tool-integrated reasoning paths, aligning with the policy and generalizing well across diverse models. Second, to perform fine-grained hierarchical optimization, we introduce an RL strategy that jointly optimizes for both episode-level problem solving and step-level code generation. This is motivated by our key insight that the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Finally, THOR incorporates a self-correction mechanism that leverages immediate tool feedback to dynamically revise erroneous reasoning paths during inference. Our approach demonstrates strong generalization across diverse models, performing effectively in both reasoning and non-reasoning models. It further achieves state-of-the-art performance for models of a similar scale on multiple mathematical benchmarks, while also delivering consistent improvements on code benchmarks. Our code will be publicly available at https://github.com/JingMog/THOR.
@@ 1 Introduction
Recent tool-free methods for enhancing LLMs’ mathematical reasoning can be broadly categorized into search-based methods (besta2024graph; zhang2025lessons; hu2025prm) and reinforcement learning (RL) paradigms (yu2025dapo; yue2025vapo; luo2026unveilingcognitivecompasstheoryofmindguided). Despite notable progress, these approaches remain constrained by a fundamental weakness of LLMs. As probabilistic next-token predictors, they inherently struggle with high-precision tasks (chen2022program), such as numerical computation, equation solving, symbolic manipulation (pan2025enhancing), and formal proofs (lewkowycz2022solving). In contrast, programmatic reasoning excels in these domains. Therefore, integrating the semantic reasoning capabilities of LLMs with the precise and verifiable execution of external code-based tools provides a promising pathway to overcome these limitations.
@@
To address these challenges, we propose THOR, a tool-integrated framework designed to enhance the reasoning ability of LLMs. (1) For constructing TIR data, in order to efficiently generate policy-aligned TIR data, we propose TIRGen, an generator-refiner-based data construction pipeline. The generator is responsible for generating natural language reasoning steps, while the refiner evaluates whether steps can be transformed into executable code and interacts with an external executor to refine the reasoning. This iterative process yields a TIR dataset that is naturally aligned with the generator’s policy and broadly applicable across diverse models and tools. (2) For performing fine-grained optimization, we are motivated by the key insight that the success of an intermediate tool call is a strong predictor of the final answer’s correctness. Our experiments later confirm this insight. Based on this, we introduce a hierarchical RL strategy that combines episode-level and step-level optimization. At the episode-level, we directly optimize for the correctness of the final answer. Concurrently, at the step-level, we apply fine-grained optimization to execution failure steps, specifically enhancing the model’s code generation ability. (3) For enhancing inference, we propose a self-correction mechanism that leverages immediate feedback from tools to dynamically revise its CoT during inference. When code invocation fails, it backtracks and explores alternative reasoning paths, thereby significantly enhancing the model’s reasoning robustness and overall performance.
@@ 2.2 TIRGen: TIR Data Generation Pipeline
Figure 2: The TIR data construction pipeline. In this pipeline, the Generator agent generates reasoning steps. The Refiner agent identifies tool-executable steps and converts them into tool-augmented reasoning steps. After multi-stage filtering, we obtain the cold start dataset $\mathcal{D}_{SFT}$.
Existing methods for TIR highlight a significant need for high-quality training data. Most approaches rely on simply prompting external large models to synthesize TIR data for non-reasoning models (gou2023tora; li2025torl), but these approaches fail to generalize effectively to reasoning models such as DeepSeek-R1. Although START (li2025start) constructs long-CoT TIR data using a rule-based prompt-hint approach, the resulting trajectories often contain redundant code invocations. Therefore, existing TIR data construction methods face critical shortcomings, including style mismatches between the generated data and policy models, as well as poor scalability to reasoning models. To overcome these limitations, we introduce TIRGen, an automated TIR data synthesis pipeline based on an generator-refiner framework, as illustrated in Figure 2.
In this framework, the Generator generates a natural language reasoning step with a maximum length of $L_{\text{step}}$. The Refiner agent then evaluates this step to identify operations that can be more efficiently solved with code, such as numerical calculations or equation solving. Upon identifying such an operation, the Refiner converts it into an executable Python code snippet while preserving the Generator’s original reasoning logic. The Refiner then interacts with a code interpreter to obtain a precise execution result, which replaces the original operation and produces a code-augmented reasoning step. The Generator resumes reasoning from this revised step, and the iterative cycle continues until a complete solution is reached, as described in Algorithm 1. This design yields two key advantages:
  • Policy Alignment. Because the Refiner observes only isolated reasoning steps without the full problem statement and final answer, the synthesized data directly reflects the Generator’s intrinsic reasoning preferences. This ensures that the data remains in-distribution, thereby mitigating the performance degradation commonly caused by training on out-of-distribution data (gudibande2023false; chen2024self).
  • Reduced Reliance on Large-scale Models. The Generator is responsible for high-level mathematical reasoning, while the Refiner requires only basic instruction-following and code-generation skills. This division of labor decomposes the complex task into subproblems that are collaboratively solved by the two agents.
@@
1:Input: Generator model $\pi_{\text{gen}}$, Refiner model $\pi_{\text{refiner}}$, Dataset $\mathcal{D}_{q}$, Code interpreter sandbox $S$.
@@
6:   $r^{t}\sim\pi_{\text{gen}}(\cdot\mid\tau),\>|r^{t}|\leq L_{\text{step}}$ $\triangleright$ Generator $\pi_{\text{gen}}$ generates a reasoning step
@@
8:     $r^{t}_{\text{logic}}\leftarrow\text{ExtractLogic}_{\pi_{\text{refiner}}}(r^{t})$ $\triangleright$ Step 1: Extract the pure reasoning part
9:     $a^{t}\leftarrow\text{ConvertToCode}_{\pi_{\text{refiner}}}(r^{t},r^{t}_{\text{logic}})$ $\triangleright$ Step 2: Convert calculation part to code
@@ 2.3.2 Hierarchical Reinforcement Learning
where $s^{\prime}_{i}$ and $A^{\prime}_{i}$ denote the responses and advantages at the step-level. In step-level optimization, each sample contains only a single think–act–observe loop, and its reward is directly determined by whether the corresponding code execution succeeds. For a group of step-level rollouts sampled from the same prefix $\{s_{i}^{\prime}\}_{i=1}^{G}$, the advantage is computed as:
$A_{i}^{\prime}=\frac{r_{i}^{\prime}-\mathrm{mean}(r^{\prime})}{\mathrm{std}(r^{\prime})}$ (9)
where $r^{\prime}$ denotes the code-execution reward within the group. The final training objective combines both episode-level and step-level losses: